# Stereo Hand-Object Reconstruction for Human-to-Robot Handover

### v2

@@ IV-B Handover setup
We follow a benchmarking protocol [4] that includes a setup of the robotic and perception systems, and 288 handover configurations to execute. The setup includes a 6-DoF Universal Robots UR5 robotic arm, equipped with a Robotiq 2F-85 2-finger gripper for grasping the object, and two Intel Realsense D435i cameras mounted on tripods on either side of the robotic arm pointing at the workspace of the robot which is a table with the delivery target marked. The protocol configurations are executed by 4 participants who hand over 3 cups and 1 plastic wine glass, either empty or filled with rice, using 3 different grasp types at 3 handover locations (72 configurations for each participant) [^3] Approval for this research was obtained from the Queen Mary Ethics of Research Committee Ref:QMERC20.525 on 19/10/2021. Consent from each human subject was obtained prior to the experiments.. The protocol instructs to initially place the object on the table at the centre of the delivery target, and then a person picks up and hands the object over to the robot. On the contrary, we assume the object to be already held by the person for our pipeline and therefore we asked participants to hold the object behind a starting line indicated by tape on the wooden table. The robot starts moving towards the object once the human moves the object beyond the starting line. To assess generalisation to different novel objects outside of the benchmark, each participant execute another 24 configurations for 8 household objects at the 3 different handover locations using 1 grasp type (a total of 96 configurations). These additional objects vary in size, topology, and appearance (Fig. 5).
Figure 5: Synthetic object types used for training StereoHO, and objects for testing hand-object reconstruction and handover experiments. For handovers: 3 cups and 1 glass, empty or filled with rice, from a benchmarking protocol [4] (left), and 8 household objects (right) to assess generalisation.
Figure 6: Comparing single-view and stereo hand object reconstructions on DexYCB [18]. For single-view, each reconstruction corresponds to the image on the same row. For stereo, the same reconstruction is shown from two viewpoints. Our method yields less noisy reconstructions and improves both the hand and object shape by combining the single view predictions.
@@ IV-C Performance measures
$\mu=\begin{cases}1-\frac{|m-\hat{m}|}{\hat{m}},&\text{if }|m-\hat{m}|<\hat{m},\\ 0,&\text{otherwise},\\ \end{cases}$ (6)
We measure grasping success as the configurations where the robot is able to grasp and hold the object without dropping it after retracting.
@@ IV-D Hand-object reconstruction from RGB
Since our reconstruction method combines two individual single-view predictions to form the stereo prediction, we evaluate our method in both single-view and stereo settings. We compare StereoHO with IHOI [20] in the single-view setting and with SVHO [27] in the stereo setting. IHOI is an SDF-based method that predicts the shape of the object conditioned on the pose of the hand, and is trained end-to-end. SVHO is a sparse multi-view method that reconstructs hand-object shape and can be easily adapted to use only two views. For a fair comparison, we train our method, IHOI, and SVHO, only on the synthetic data of ObMan, and evaluate the methods using DexYCB [18]. We select the two views closest to the stereo views of our setup and compare the reconstruction quality (Fig. 6). For the single-view setting, we aggregate results from both the left and right views. We separate the testing set into seen and unseen categories based on the similarity in 3D shape between the training and testing objects (Fig. 5).
@@
Figure 7: Single-view and stereo hand-object reconstruction errors on DexYCB [18]. Best results on the bottom-left corner. Legend: CD: chamfer distance,  SVHO [27],  IHOI [20],  Ours (no segmentation mask),  Ours. IHOI, as a single-view reconstruction method, is not evaluated in stereo setting.
@@
In the single-view setting (Fig. 7), our method outperforms IHOI on object reconstruction by 9.71 cm2 and 12.52 cm2, while slightly underperforming on hand reconstruction by 1.10 cm2 and 1.21 cm2 in the seen and unseen categories, respectively. In the stereo setting, our method outperforms SVHO on object reconstruction by 35.43 cm2 and 32.27 cm2, while performing slightly better on the hand reconstruction by 0.51 cm2 and 0.13 cm2 in the seen and unseen categories, respectively. Compared with a version of our method without segmentation masks as input, our method improves the object reconstruction while sacrificing a small amount of performance in hand reconstruction. This also results in less noisy reconstructions as shown in Fig. 6 when comparing with IHOI in the single-view setting.
@@ IV-E Human-to-robot handovers
We compare StereoHO with a depth-based baseline (DB), ClearGrasp [11], and the CORSMAL baseline (CB) [6]. DB uses the pointcloud observed from the left camera. ClearGrasp is a depth completion method to handle transparent objects using depth-based sensing. For fairness, DB and ClearGrasp are integrated in the same handover pipeline. CB [6] is a stereo-based pipeline specifically designed for handovers of rotationally symmetric container objects. Unlike our pipeline, CB assumes the object is initially placed on the table, following the original protocol [4], and is upright throughout the entire handover process. The object is detected using Mask R-CNN [39], pretrained on COCO [40], and the object shape is estimated with a multi-view iterative fitting algorithm given a cylindrical prior [17]. For each frame, hand keypoints are estimated using OpenPose [41] and the 3D object centroid is triangulated from the 2D centroid in the left and right views after tracking the object mask in each view with SiamMask [42]. Grasp points are generated along the height of the container and those ones close to the hand keypoints on the vertical axis are removed. The robot starts moving once the object is lifted off the table.
@@
Tab. I compares the handover performance of the methods. For the container objects from the CORSMAL benchmark [4], both CB and our method achieve over 75% grasping success and 65% delivery success. This shows that assuming a cylindrical prior and a mostly upright pose throughout the task is reasonable when handing over cups and drinking glasses [6]. When handing over general household objects, our method achieved 75% grasping success on the opaque objects and 91% grasping success on the transparent containers, significantly outperforming DB and ClearGrasp. Comparing the hand and object shape reconstructed by each method (Fig. 8), DB is easily affected by transparency of the object and noisy depth measurements, leaving holes in the reconstructed shape. ClearGrasp can fill in the missing depth value but the estimated value is often inaccurate and noisy. Our method produces a complete shape which helps the grasp estimation method to produce successful grasp candidates for the handover.
Figure 9: Examples of successful human-to-robot handovers of household objects using StereoHO, and failure cases due to grasping failure or spilled filling.
Overall, we showed that our method can perform human-to-robot handovers on general household objects (Fig. 9), outperforming depth-based methods and performing on par on containers with a method designed specifically for the handover of containers. Specifically, our method can successfully receive objects with challenging shapes, such as thin (screwdriver), deformable (cloth), and reflective objects (CD), whereas previous stereo-based method fail to generalise to non-container objects, and depth-based methods fail to receive transparent objects. Failures occur when the reconstruction of the object shape is inaccurate. For example, the robot often mistakenly try to grasp the cracker box from the wider side (Fig. 9), which is larger than the gripper width. Failure can also occur when the robot loses track of the hand due to occlusions by the object or the robot arm itself.
@@ V Conclusion
In this paper, we proposed a joint hand-object reconstruction method to enable human-to-robot handovers using wide baseline stereo RGB cameras. Our reconstruction method outperforms existing methods in both single-view and stereo settings. Handover experiments showed that our method enables a robot to receive empty and filled containers, and can generalize to other household objects. Future work includes improving the speed of handovers and the quality of hand-object reconstructions.
ACKNOWLEDGMENT
This project made use of time on Tier 2 HPC facility JADE2, funded by EPSRC (EP/T022205/1).

### v3

@@ IV-B Handover setup
We follow a benchmarking protocol [4] that includes a setup of the robotic and perception systems and 288 handover configurations to execute. The setup includes a 6-DoF Universal Robots UR5 robotic arm, equipped with a Robotiq 2F-85 2-finger gripper for grasping the object, and two Intel Realsense D435i cameras mounted on tripods on each side of the robotic arm pointing at the workspace of the robot which is a table with the delivery target marked. The protocol configurations are executed by 4 participants who hand over 3 cups and 1 plastic wine glass, either empty or filled with rice, using 3 different grasp types at 3 handover locations (72 configurations for each participant) [^3] Approval for this research was obtained from the Queen Mary Ethics of Research Committee Ref:QMERC20.525 on 19/10/2021. Consent from each human subject was obtained prior to the experiments.. The protocol instructs to initially place the object on the table at the centre of the delivery target, and then a person picks up and hands the object over to the robot. We assume the object to be already held by the person for our pipeline and therefore we asked participants to hold the object behind a starting line indicated by tape on the table. The robot starts moving towards the object once the human moves the object beyond the starting line. To assess generalisation to different novel objects outside of the benchmark, each participant executes 24 configurations for 8 household objects at the 3 different handover locations using 1 grasp type (a total of 96 configurations). These additional objects vary in size, topology, and appearance (see Fig. 5).
Figure 5: Synthetic object types used for training StereoHO (top row), and objects for testing hand-object reconstruction (middle row). For handovers (bottom row): 3 cups and 1 glass, empty or filled with rice, from a benchmarking protocol [4] (left) and 8 household objects (right) to assess generalisation.
Figure 6: Comparing single-view and stereo hand object reconstructions on DexYCB [18]. For single-view, each reconstruction corresponds to the image on the same row. For stereo, the same reconstruction is shown from two viewpoints. Our method yields less noisy reconstructions and improves both the hand and object shape by combining the single-view predictions.
@@ IV-C Performance measures
$\mu=\begin{cases}1-\frac{|m-\hat{m}|}{\hat{m}},&\text{if }|m-\hat{m}|<\hat{m},\\ 0,&\text{otherwise}.\\ \end{cases}$ (6)
We measure grasping success as the percentage of handover configurations where the robot grasps and holds the object without dropping it after retracting.
@@ IV-D Hand-object reconstruction from RGB
Since our reconstruction method combines two individual single-view predictions to form the stereo prediction, we evaluate our method in both single-view and stereo settings. We compare StereoHO with IHOI [20] in the single-view setting and with SVHO [27] in the stereo setting. IHOI is an SDF-based method that predicts the shape of the object conditioned on the pose of the hand, and is trained end-to-end. SVHO is a sparse multi-view method that reconstructs hand-object shape and can be easily adapted to use only two views. For a fair comparison, we train our method StereoHO, IHOI, and SVHO, only on the synthetic data of ObMan, and evaluate the methods using DexYCB [18]. We select the two views closest to the stereo views of our setup and compare the reconstruction quality (see Fig. 6). For the single-view setting, we aggregate results from both the left and right view. We separate the testing set into seen and unseen categories based on the similarity in 3D shape between the training and testing objects (see Fig. 5).
@@
Figure 7: Single-view and stereo hand-object reconstruction errors on DexYCB [18]. Best results on the bottom-left corner. Legend – CD: Chamfer distance,  SVHO [27],  IHOI [20],  Ours (no segmentation mask),  Ours. IHOI (single-view reconstruction method) is not evaluated in stereo setting.
@@
In the single-view setting (see Fig. 7), our method outperforms IHOI on object reconstruction by 9.71 cm2 and 12.52 cm2, while slightly underperforming on hand reconstruction by 1.10 cm2 and 1.21 cm2 in the seen and unseen categories, respectively. In the stereo setting, our method outperforms SVHO on object reconstruction by 35.43 cm2 and 32.27 cm2, while performing slightly better on the hand reconstruction by 0.51 cm2 and 0.13 cm2 in the seen and unseen categories, respectively. Compared with a version of our method without segmentation masks as input, our method improves the object reconstruction while sacrificing a small amount of performance in hand reconstruction. This also results in less noisy reconstructions when comparing with IHOI in the single-view setting (see Fig. 6).
@@ IV-E Human-to-robot handovers
We compare StereoHO with a depth-based baseline (DB), ClearGrasp [11], and the CORSMAL baseline (CB) [6]. DB uses the pointcloud observed from the left camera. ClearGrasp is a depth completion method to handle transparent objects using depth-based sensing. For fairness, DB and ClearGrasp are integrated in the same handover pipeline. CB [6] is a stereo-based pipeline specifically designed for handovers of rotationally symmetric container objects. Unlike our pipeline, CB assumes the object is initially placed on the table, following the original protocol [4], and is upright throughout the entire handover process. The object is detected using Mask R-CNN [39], pretrained on COCO [40], and the object shape is estimated with a multi-view iterative fitting algorithm given a cylindrical prior [17]. For each frame, hand keypoints are estimated using OpenPose [41] and the 3D object centroid is triangulated from the 2D centroids of the object masks tracked with SiamMask [42]. Grasp points are generated along the height of the container and those ones close to the hand keypoints are removed. The robot starts moving once the object is lifted off the table.
@@
Tab. I compares the handover performance of the methods. For the containers from the CORSMAL benchmark [4], both CB and our method achieve over 75% grasping success and 65% delivery success. This shows that assuming a cylindrical prior and an upright pose throughout the task is reasonable when handing over containers [6]. Note that objects that are small relative to the hand (e.g., the small white cup) are less common in the training data of StereoHO and hence are more challenging to reconstruct and grasp. Moreover, the added components compared to CB, such as the reconstruction and grasp estimation modules, contribute to a higher computational load, lowering the efficiency score. When handing over household objects, our method achieves 75% grasping success on the opaque objects and 91% grasping success on the transparent containers, significantly outperforming DB and ClearGrasp. Comparing the hand and object shape reconstructed by each method (see Fig. 8), DB is affected by the transparency of the object and noisy depth measurements, leaving holes in the reconstructed shape. ClearGrasp fills in the missing depth value. However, the estimated value is often inaccurate and noisy. StereoHO reconstructs a complete shape that helps estimate successful grasp candidates for the handover.
Figure 9: Examples of successful human-to-robot handovers of household objects using StereoHO, grasping failures, and spilling of the content.
Overall, our method outperforms depth-based methods when handing over general household objects (see Fig. 9). For the handover of containers, our method performs on par with CB, which is specifically designed for the handover of containers. Specifically, our method successfully receives objects with challenging properties, such as thin shape (screwdriver), deformable shape (cloth), and reflective material (CD). On the contrary, CB fails to generalise to non-container objects, and depth-based methods (DB and ClearGrasp) fail to receive transparent objects. Failures occur for StereoHO when the reconstruction of the object shape is inaccurate. For example, the robot often try to grasp the cracker box from the wider side (see Fig. 9), which is larger than the gripper width. Failure can also occur when the robot loses track of the hand due to occlusions by the object or the robot arm itself.
@@ V Conclusion
In this paper, we proposed a joint hand-object reconstruction method to enable human-to-robot handovers using wide baseline stereo RGB cameras. Our reconstruction method outperforms IHOI and SVHO in both single-view and stereo settings. Handover experiments showed that our method enables a robot to receive empty and filled containers, and can generalize to other household objects. Future work includes improving the speed of handovers and the quality of hand-object reconstructions.